# An International Agreement to Prevent the Premature Creation of Artificial Superintelligence

### v1

@@ Appendix A The Agreement
To help verify that there is no prohibited AI research happening, Article IX tasks parties with demarcating “areas adjacent to Restricted Research” and then establishing relationships with the researchers working in these adjacent areas. There are sufficiently few top AI researchers in the world that it may be feasible to track the activities of a significant fraction of them. The technical staff of top AI companies numbers on the order of 5,000 researchers [^21] OpenAI’s GPT-5 announcement lists 470 contributors. The Google’s Gemini 1.5 technical report lists 1135 contributors. The Llama 3 paper lists 559 authors. The DeepSeek-V3 paper lists 199 authors. The Qwen3 paper lists 177 total contributors. The Kimi k2 paper lists 168 authors. The sum of these numbers is 2,708. Other frontier AI companies such as Anthropic and xAI do not provide sufficient details about the number of technical contributors, to our knowledge. There are other companies that may have sufficient AI talent, therefore we conservatively estimate that the total number of researchers at top AI companies is on the order of 5,000., and it is commonly believed that a much smaller group is critical to frontier AI development, likely numbering in the hundreds. [^22] In a 2025 interview, David Luan, head of Amazon’s AGI research lab, estimated the number of people he would trust “with a giant dollar amount of compute” to develop a frontier model at “sub-150.” The number of attendees of top AI conferences is estimated to be about 70,000. As higher-end estimates, the number of employees at relevant hardware companies likely numbers around a million [^23] TSMC employee count: 84,000 in 2025. Intel employee count: 125,000 in 2023. Samsung employee count: 263,000 in 2024. SMIC employee count: 20,000 in 2023. NVIDIA employee count: 36,000 in 2025. AMD employee count: 26,000 in 2023. Huawei employee count: 208,000 in 2024. Google TPU division: unknown. ASML employee count: 44,000 in 2024. SK Hynix employee count: 47,000 in 2024. Micron employee count: 43,000 in 2023. The sum of these numbers is 896,000. So the total number of employees at the most relevant AI hardware companies is likely around one million, though the core technical staff are probably a small subset of this and the number of crucial researchers may only be a few thousand. and the number of people with at least basic technical knowledge of AI knowledge is likely in the single digit or tens of millions [^24] In April 2025, the popular platform for sharing AI models and datasets HuggingFace had 8 million users. The number of students who have taken Andrew Ng’s Machine Learning course is about 4.8 million. The number of developers using github (broader than just AI) was around 100 million in 2023, and the total number of professional software developers in the world is estimated at 47 million in 2025.. States could interview researchers about their activities and offer asylum and financial incentives for any whistleblowers (see Article X).
@@
      Parties prohibit and prevent retaliation against Covered Whistleblowers and Associated Persons, including but not limited to dismissal, demotion, blacklisting, loss of benefits, harassment, intimidation, threats, civil or criminal actions, visa cancellation, physical violence, imprisonment, restriction of movement, or other adverse measures. Any contractual terms (including non-disclosure or non-disparagement agreements) purporting to limit protected disclosures under this Agreement are void and unenforceable. Mistreatment of whistleblowers constitutes a violation of this Agreement and is handled under Article XI, paragraph 3.
@@
      Parties offer asylum or humanitarian protection to Covered Whistleblowers and their families, provide safe-conduct travel documents, and coordinate secure transit.
@@
    Recognizing that the development of ASI or other Dangerous AI Activities, as laid out in Articles IV through IX, would pose a threat to global security and to the life of all people, it may be necessary for Parties to this Agreement to take drastic actions to prevent such development. The Parties recognize that development of artificial superintelligence (ASI), anywhere on earth, would be a threat to all Parties. Under Article 51 of the United Nations Charter and as longstanding precedent, states have a right to self-defence. Due to the scale and speed of ASI-related threats, self-defence may require pre-emptive actions to prevent the development of ASI.
@@
    To prevent the development or deployment of ASI, this Article authorizes tailored Protective Actions. Where there is credible evidence that a State or other actor (whether a Party or a non-Party) is conducting or imminently intends to conduct activities aimed at developing or deploying ASI in violation of Article I, Article IV, Article V, Article VI, Article VII, or Article VIII, a State Party may undertake Protective Actions that are necessary and proportionate to prevent activities. In recognition of the harms and escalatory nature of Protective Actions, Protective Actions should be used as a last-resort. Outside of emergencies and time-sensitive situations, Protective Actions are preceded by other approaches such as, but not to limited to:
@@
      Identifies the specific AI-enabling activities and assets targeted;
@@
      Verified surrender or destruction of covered chip clusters or ASI-enabling assets, potentially including the establishment of sufficient safeguards to prevent Restricted Research activities.
@@
    For AI models created via declared training or post-training within the limits of Article IV, the CTB may require evaluations and other tests. These tests will inform whether the thresholds set in Article IV, Article V, Article VII, and Article VIII need to be revised. The methods used for reviews will be determined by the CTB and may be updated.
@@ Appendix B What can we do today?
          Work with chip companies to enhance chip security and lay the ground for future verification measures.
@@
    • Fund research focused on AI model evaluations, safety research measures to mitigate risk, and trigger points for implementation of such measures.
    • Invest in AI verification R&D, hardware-level controls, and other technology development that would aid future governance.
@@
    • Enhance security for current and future AI projects, in order to protect against proliferation to and misuse by third parties, such as terrorists or rogue states.
@@
    • Develop incident response capabilities [somani2025strengthening].
@@ Appendix C Stages
International discussions on AI governance and the establishment of guardrails remain at a nascent stage. Predicting the path such negotiations might take is inherently difficult, but this section illustrates one possible pathway to governance—one that begins with bilateral measures between the U.S. and PRC and gradually expands to involve relevant states.
@@
Stage Two – Building Transparency and Communication

### v2

@@ Appendix A The Agreement
To help verify that there is no prohibited AI research happening, Article IX tasks parties with demarcating “areas adjacent to Restricted Research” and then establishing relationships with the researchers working in these adjacent areas. There are sufficiently few top AI researchers in the world that it may be feasible to track the activities of a significant fraction of them. The technical staff of top AI companies numbers on the order of 5,000 researchers [^20] OpenAI’s GPT-5 announcement lists 470 contributors. The Google’s Gemini 1.5 technical report lists 1135 contributors. The Llama 3 paper lists 559 authors. The DeepSeek-V3 paper lists 199 authors. The Qwen3 paper lists 177 total contributors. The Kimi k2 paper lists 168 authors. The sum of these numbers is 2,708. Other frontier AI companies such as Anthropic and xAI do not provide sufficient details about the number of technical contributors, to our knowledge. There are other companies that may have sufficient AI talent, therefore we conservatively estimate that the total number of researchers at top AI companies is on the order of 5,000., and it is commonly believed that a much smaller group is critical to frontier AI development, likely numbering in the hundreds. [^21] In a 2025 interview, David Luan, head of Amazon’s AGI research lab, estimated the number of people he would trust “with a giant dollar amount of compute” to develop a frontier model at “sub-150.” The number of attendees of top AI conferences is estimated to be about 70,000. As higher-end estimates, the number of employees at relevant hardware companies likely numbers around a million [^22] TSMC employee count: 84,000 in 2025. Intel employee count: 125,000 in 2023. Samsung employee count: 263,000 in 2024. SMIC employee count: 20,000 in 2023. NVIDIA employee count: 36,000 in 2025. AMD employee count: 26,000 in 2023. Huawei employee count: 208,000 in 2024. Google TPU division: unknown. ASML employee count: 44,000 in 2024. SK Hynix employee count: 47,000 in 2024. Micron employee count: 43,000 in 2023. The sum of these numbers is 896,000. So the total number of employees at the most relevant AI hardware companies is likely around one million, though the core technical staff are probably a small subset of this and the number of crucial researchers may only be a few thousand. and the number of people with at least basic technical AI knowledge is likely in the single digit or tens of millions [^23] In April 2025, the popular platform for sharing AI models and datasets HuggingFace had 8 million users. The number of students who have taken Andrew Ng’s Machine Learning course is about 4.8 million. The number of developers using github (broader than just AI) was around 100 million in 2023, and the total number of professional software developers in the world is estimated at 47 million in 2025.. States could interview researchers about their activities and offer asylum and financial incentives for any whistleblowers (see Article X).
@@
      Parties prohibit and prevent retaliation against Covered Whistleblowers and Associated Persons, including but not limited to dismissal, demotion, blacklisting, loss of benefits, harassment, intimidation, threats, civil or criminal actions, visa cancellation, physical violence, imprisonment, restriction of movement, or other adverse measures. Any contractual terms (including non‑disclosure or non‑disparagement agreements) purporting to limit protected disclosures under this Agreement are void and unenforceable. Mistreatment of whistleblowers constitutes a violation of this Agreement and is handled under Article XI, paragraph 3.
@@
      Parties offer asylum or humanitarian protection to Covered Whistleblowers and their families, provide safe‑conduct travel documents, and coordinate secure transit.
@@
    Recognizing that the development of ASI or other Dangerous AI Activities, as laid out in Articles IV through IX, would pose a threat to global security and to the life of all people, it may be necessary for Parties to this Agreement to take drastic actions to prevent such development. The Parties recognize that development of artificial superintelligence (ASI), anywhere on earth, would be a threat to all Parties. Under Article 51 of the United Nations Charter and as longstanding precedent, states have a right to self-defense. Due to the scale and speed of ASI-related threats, self-defense may require pre-emptive actions to prevent the development of ASI.
@@
    To prevent the development or deployment of ASI, this Article authorizes tailored Protective Actions. Where there is credible evidence that a State or other actor (whether a Party or a non‑Party) is conducting or imminently intends to conduct activities aimed at developing or deploying ASI in violation of Article I, Article IV, Article V, Article VI, Article VII, or Article VIII, a State Party may undertake Protective Actions that are necessary and proportionate to prevent activities. In recognition of the harms and escalatory nature of Protective Actions, Protective Actions should be used as a last-resort. Outside of emergencies and time-sensitive situations, Protective Actions are preceded by other approaches such as, but not limited to:
@@
      Identifies the specific AI‑enabling activities and assets targeted;
@@
      Verified surrender or destruction of covered chip clusters or ASI‑enabling assets, potentially including the establishment of sufficient safeguards to prevent Restricted Research activities.
@@
    For AI models created via declared training or post‑training within the limits of Article IV, the CTB may require evaluations and other tests. These tests will inform whether the thresholds set in Article IV, Article V, Article VII, and Article VIII need to be revised. The methods used for reviews will be determined by the CTB and may be updated.
@@ Appendix B What can we do today?
          Conduct R&D for on-chip hardware-enabled governance mechanisms (HEMs) and for supporting equipment that could verify the location and the types of workloads conducted on advanced AI chips (e.g., training, inference). Develop tamper-resistant HEM technologies and explore advanced HEMs capabilities that allow for remote governance, licensing, and limits on chip use.
          Establish an interagency group within the US government to rapidly develop the framework for implementing location verification and advanced HEMs.
          Establish a public/private partnership to develop advanced HEM technologies
@@
    • Fund research focused on AI evaluations (both before and after model deployment), safety research measures to mitigate risk, and trigger points for implementation of such measures.
    • Invest in verification expertise and know-how, for instance by funding pilot verification efforts using open source intelligence or satellite data.
@@
    • Enhance security for current and future AI projects, in order to protect against proliferation to and misuse by third parties including terrorists or rogue states.
@@
    • Task an executive branch agency with assessing the implications of advanced AI on key national security areas including CBRN and cyber capabilities, model autonomy, and automated AI R&D to speed up AI developments, and open-source models.
    • Develop scenario-based incident response capabilities [somani2025strengthening].
@@ Appendix C Stages
International discussions on AI governance and the establishment of guardrails remain at a nascent stage. Predicting the path such negotiations might take is inherently difficult, but this section illustrates one possible pathway to governance—one that begins with bilateral measures between the U.S. and PRC and gradually expands to involve relevant states. This should be considered a hypothetical scenario rather than a recommendation.
@@
Stage Two – Enhanced Transparency and Communication